### v1

Abstract
Divisor methods are well known to satisfy house monotonicity, which allows representative seats to be allocated sequentially. We focus on stationary divisor methods defined by a rounding cut point $c\in[0,1]$. For such methods with integer-valued votes, the resulting apportionment sequences are periodic. Restricting attention to two-party allocations, we characterize the set of possible sequences and establish a connection between the lexicographical ordering of these sequences and the parameter $c$. We then show how sequences for all pairs of parties can be systematically extended to the $n$-party setting. Further, we determine the number of distinct sequences in the $n$-party problem for all $c$. Our approach offers a refined perspective on large-party bias: rather than viewing large parties as simply receiving more seats, we show that they instead obtain their seats earlier in the apportionment sequence. Of particular interest is a new relationship we uncover between the sequences generated by the smallest divisors (Adams) and greatest divisors (d’Hondt or Jefferson) methods.
keywords:
apportionment sequence, divisor methods, d’Hondt/Jefferson method, Adams method
pacs:
MSC Classification]91B32 91B12 91F10
1 Introduction
Divisor methods of apportionment are widely used in proportional representation systems to allocate seats to political parties in proportion to their share of the popular vote. Less attention, however, has been given to how these methods determine the sequence in which seats are awarded. This sequence is important in practice as it is used by some coalition governments to determine cabinet positions: while the office of prime minister typically goes to the party with a plurality of votes, other cabinet positions are selected by parties based on the sequence. For example, in Northern Ireland, the greatest divisors method (also known as d’Hondt’s or Jefferson’s method) is used not only to determine the number of seats allocated to each party in the Assembly but also the order in which cabinet posts are selected [14]. Sequential apportionment in this context offers an alternative to the intense negotiation and strategic bargaining that often arise during the establishment of a new cabinet. Using an apportionment method to allocate positions sequentially can help mitigate these conflicts and avoid delays.
Just-in-Time (JIT) sequencing problems have the same structure as the sequential allocation of minister positions. In these problems, a number of items must be processed (or produced) to meet a set of demands or users. The items must be processed one at a time, and so the question arises as to what item order satisfies the demands in an equitable way. Although divisor methods have been used to allocate the processing of items, other algorithms have been developed to solve such problems under different optimality conditions, such as minimizing the total deviation or the maximum deviation. Kubiak [12] provides a comprehensive analysis of JIT sequencing problems, exploring the usefulness of divisor methods for JIT sequencing and analyzing the complexity of algorithms used to solve JIT sequencing problems.
Our work has application to sequential apportionment in both fields, as we focus our attention on the possible apportionment sequences from stationary divisor methods, a family parametrized by a rounding cut point $c\in[0,1]$ which includes the smallest divisors method (also known as Adams’ method for which $c=0$), Webster’s or Sainte Laguë’s method ($c=0.5$), and d’Hondt’s method (used in Northern Ireland, for which $c=1$). Such sequences arise whenever apportionment methods satisfy the property of house monotonicity, which requires that no party’s allocation decreases as the number of total available seats increases. Monotonic apportionment methods give rise naturally to a sequence by considering the allocation of one seat (the first seat), the allocation of the second seat, and so on. Stationary divisor methods have an easily described algorithm for determining this sequence—versions of which are used in awarding minister positions and in JIT sequencing problems.
Stationary divisor methods also have a nice property in that the sequences they give rise to are periodic (when vote totals are integers), as proved by Kubiak [12] in the context of JIT sequencing. Not all divisor methods yield periodic sequences, as evidenced by Hill-Huntington’s method, which is used in the U.S. to determine the numbers of seats each state receives in the U.S. House of Representatives. Periodicity implies that there are only a finite number of possible sequences from stationary methods. In this paper, we identify all possible sequences generated by stationary divisor methods for given party vote totals and provide explicit formulas for the total number of such sequences as functions of party vote shares. This is in contrast to work by Cembrano et al. [4], in which they use a geometric approach to obtain bounds on the number of different possible apportionments (not sequential apportionments) for $n$ parties for the class of stationary divisor methods, as well as for power-mean divisor methods. We also establish necessary and sufficient conditions for determining when a periodic sequence corresponds to a sequential apportionment, as well as methods for identifying the associated stationary divisor method.
We combine the periodicity of the apportionment sequences and a notion of weak proportionality for divisor methods to refine the notion of size bias of divisor methods. As identified in Balinski and Young [2], Adams’ method provides one extreme that favors small parties and the d’Hondt method provides the other extreme that favors large parties. However, by applying a lexicographic ordering of apportionment sequences, we reframe this distinction as less about an absolute bias and more about variation in when large and small parties receive their seats. The refinement also leads to a new relationship between both the sequences and apportionments from Adams’ and d’Hondt’s methods.
Applying a sequential framework to the apportionment problem is a relatively unstudied field. In addition to the literature analyzing political apportionment cited above, Pukelshein [15] provides an in-depth study of the use of divisor methods in systems of proportional representation, but does not focus on sequential apportionment. A small number of studies have examined the use or potential use of divisor sequences for the distribution of ministerial positions. O’Leary, Grofman, and Elkit [14] analyze the systems used in the Northern Ireland Assembly, in Danish municipal governments, and in the European parliament to determine committee chairs. Raabe and Linhart [16] investigate several models incorporating party preferences and the sequential allocations arising from the d’Hondt, the Sainte Laguë, and Adams’ methods and compare them to actual cabinet allocations in the German parliament. They note that adapting the structure of such sequences may be an advantage to parties given the time pressures they face in selecting cabinet portfolios after an election. Ecker, Meyer, and Müller [8] take a similar modeling approach in analyzing party portfolios for 146 coalition governments in Western and Central Eastern Europe.
When cabinet ministry positions are awarded sequentially, Brams and Kaplan [3] explain that it may not be rational for parties to naively select their most-preferred minister when available. Further, they show that strategic selection may result in a combination of ministry positions that is not Pareto-optimal, making all parties worse off, and how choosing earlier in the sequence may be harmful to a party. For two parties, Brams and Kaplan [3] show how to combine sequential choices with a type of trading to eliminate these negative consequences.
The literature on sequential apportionment in JIT problems is more developed. Dhamala and Thapa [6] provide an overview of the relationship between JIT and apportionment problems, while Dhamala, Thapa and Yu [7] propose new mean-based divisor methods for minimizing the sum of deviations in a JIT sequencing problem. More along the approaches taken in social choice theory, Józefowska, Józefowsk, and Kubiak [10] provide an axiomatic approach to JIT sequencing problems, while Chun [5] focuses on fair queueing and strategic interactions. Kubiak [12] shows that fair queueing based on starting times results in the sequence from Adams’ method, while fair queueing based on finishing times results in the sequence from d’Hondt’s method. Józefowska, Józefowsk, and Kubiak [11] show that no divisor method solves the Liu-Layland periodic scheduling problem, which is similar to JIT sequencing problem except it includes times by which jobs must be finished.
In the next section, we review the basics of divisor methods and their application to award representatives in succession; this includes properties satisfied by divisor methods, but also new results about apportionment sequences and their period length. Section 2 also explains the partial ordering on apportionment sequences and concludes with a result tying the apportionment sequences from the divisor methods of Adams and d’Hondt.
In Section 3, we focus our attention on properties of 2-party apportionment sequences. For any vote distribution, we characterize the possible 2-party apportionment sequences for all stationary divisor methods, breaking the interval of possible cut points $[0,1]$ into equivalence classes. The possible apportionment sequences for $n$ parties can be determined by lifting all $\binom{n}{2}$ pairs of 2-party sequences to the $n$-party sequence. Using properties of the common divisors of the parties’ vote totals, we are able to count the number of possible sequences for the $n$-party problem, discussed in Section 4. We offer some concluding remarks in Section 5.
2 Sequential Apportionment
Proportional representation systems assign seats to parties in roughly the same proportion as the votes they receive. The d’Hondt method—also called the greatest divisors method—assigns seats to each party sequentially by dividing their vote count by successive divisors $1,2,3,\dots$ and allocating the next seat to the party with the greatest ratio. At each step, the divisor for the party that received the previous seat is updated. The process is repeated until all seats have been allocated. We illustrate with an example from the 2022 Northern Ireland Assembly. This election was notable in part because it was the first election in which Sinn Féin was awarded the largest number of seats. Eight parties received seats, based on the number of votes received. Following the election, positions for the Executive Committee were selected sequentially using the d’Hondt method. However, instead of using the popular vote totals, the proportions were based on the number of seats received by each party in the Assembly.
Example 1.
The 2022 election for the Northern Ireland Assembly resulted in the following eight parties receiving seats:
Sinn Féin (27) Democratic Unionist Party (25)
Alliance Party (17) Ulster Unionist Party (9)
Social Democratic & Labour Party (8) Independent (2)
Traditional Unionist Voice (1) People Before Profit (1)
The top two vote-getting parties, Sinn Féin (SF) and the Democratic Unionist Party (DUP), nominated individuals for the First Minister and the deputy First Minister, respectively. The position of Justice Minister was awarded to the Alliance Party (AP) as part of a political compromise requiring cross-community support from both nationalists and unionists. The remaining positions were then allocated to the parties sequentially. For instance, the next position was allocated to SF because
$\frac{27}{1}=\max\left\{\frac{27}{1},\frac{25}{1},\frac{17}{2},\frac{9}{1},\frac{8}{1},\frac{2}{1},\frac{1}{1}\right\},$
where the denominator for the Alliance Party (AP) was increased to 2 to reflect their being awarded the Justice Minister position. The next position was given to the DUP because
$\frac{25}{1}=\max\left\{\frac{27}{2},\frac{25}{1},\frac{17}{2},\frac{9}{1},\frac{8}{1},\frac{2}{1},\frac{1}{1}\right\},$
because the divisor for SF had been updated in a similar manner.
Continuing in this way, the positions were eventually allocated as follows: Departments of the Economy (SF), of Education (DUP), of Finance (SF), for Communities (DUP), of Health (UUP), of Infrastructure (SF), and of Agriculture, Environment and Rural Affairs (AP). This is a result of the following set of inequalities
SF DUP SF DUP UUP SF AP
$27/1$ >> $25/1$ >> $27/2$ >> $25/2$ >> $9/1$ $\geq$ $27/3$ >> $17/2$
Even though the use of sequential apportionment is supposed to mitigate delays in the negotiation of minister positions, the Democratic Unionist Party still delayed the formation of the Executive Committee, which resulted in legal action to prevent a snap election. Thus, this selection process was not completed until 2024, and the use of sequential apportionment did not in this case decrease the delay in forming the Executive Committee.
The use of the divisors $1,2,3,\ldots,$ distinguishes D’Hondt’s method. More generally, divisor methods can be defined through their unique set of denominators or divisors. In this article, we focus on the class of stationary divisor methods, which are parametrized by a fixed constant $c\in[0,1]$, and given by the sequence $a+c$ where $a$ is the nonnegative integer number of positions already awarded, that is, the current apportionment. We refer to the value $c$ as the cut point of the corresponding divisor method. The process of sequential apportionment for a general cut point $c$ is defined below.
Let $p_{i}\in\mathbb{N}$ be the number of votes that party $i$ receives for $i\in N=\{1,2,\dots,n\}$ and $h\in\mathbb{N}$ be the number of seats to be awarded, usually referred to as the house size. Without loss of generality, we assume that parties are ordered so that $p_{i}\geq p_{j}$ if $i>j$.
Definition 1.
For fixed $0\leq c\leq 1$, given a distribution of votes $\mathbf{p}=(p_{1},p_{2},\ldots,p_{n})$, the apportionment sequence for the stationary divisor method with cut point $c$ is defined inductively as follows. Start by awarding each party $a_{i}=0$ seats. Then if party $p_{i}$ has been awarded $a_{i}$ seats, award the next seat to the party that maximizes the ratio $\frac{p_{i}}{a_{i}+c}$. Continue until all $h$ seats have been allocated.
If $c=1$, then the sequential apportionment process corresponds to the D’Hondt method. If $c=1/2$, the process corresponds to the Sainte-Lagué or Webster method. If $c=0$, the apportionment method is known as Adams’ method. Because $a_{i}=0$ for all parties $i$ before any seats have been allocated, Adams’ method is undefined because $\frac{p_{i}}{a_{i}}$ is undefined, because dividing by zero is undefined. To apply Adams’ method, it is assumed that $\frac{p_{i}}{0}>\frac{p_{j}}{0}$ whenever $p_{i}>p_{j}$ and $\frac{p_{i}}{0}>\frac{p_{j}}{k}$ for any $k>0$. Thus, under Adams’ method, the first $n$ seats are awarded, one to each party, assuming there are sufficient seats. There are other divisor methods, such as the Hill-Huntington method and Dean’s method, based on the geometric and harmonic means, respectively, that also give a seat to every party before a party gets a second seat. (To see how Hill-Huntington’s method fits into the sequential framework, it awards the next seat to the party that maximizes the ratio $p_{i}/\sqrt{a_{i}(a_{i}+1)}$, where $\sqrt{a_{i}(a_{i}+1)}$ is the geometric mean of the current apportionment $a_{i}$ and the next apportionment $a_{i}+1$.)
Because of the possibility of ties, the apportionment sequence may not be unique. We assume that ties are broken in favor of the larger parties. Therefore, if at some stage in the apportionment sequence, $p_{i}/(a_{i}+c)=p_{j}/(a_{j}+c)=\max_{k\leq n}p_{k}/(a_{k}+c)$, then the seat is awarded to the party with the smallest index $i$, that is, the party with the most votes, unless there were parties with the same largest number of votes.
In practice, vote totals are very large. However, divisor methods depend only on the ratios between vote distributions. Hence, our examples will use relatively small $p_{i}$ values.
Example 2.
Suppose two parties receive votes $\mathbf{p}=(32,14)$ with a house size of $h=5$. If $c=1$ then the first two seats are given to party 1 since $32/1>32/2>14/1$. The third and fourth seats are given to party 2 and party 1, respectively, since $14/1>32/3>14/2$. The fifth seat is given to party 2 since $32/4>14/2$.
If $c=0$, then each party receives a seat: party 1 receives the first seat and party 2 receives the second seat. The third and fourth seats are given to $1$ since, again, $32/1>32/2>14/1$. The fifth seat is given to $2$ since $\frac{14}{1}>\frac{32}{3}$.
In the example, notice that after 5 seats have been given out, parties 1 and 2 have received 3 and 2 seats, respectively, under both D’Hondt’s and Adams’ methods, however the sequence with which these seats have been allocated is different. We focus on the apportionment sequences that include apportionments for all house sizes $h$ by considering the infinite sequence.
Definition 2.
For fixed $\mathbf{p}$, cut point $c$, and a house size $h$, let $S_{c}(h,\mathbf{p})=(s_{1},s_{2},\dots,s_{h})$ where $s_{j}=i$ if party $i$ receives the $j$th seat for each $j=1,\ldots,h$. Let $S_{c}(\mathbf{p})=\lim_{h\rightarrow\infty}S_{c}(h,\mathbf{p})$.
Thus, in Example 2, $S_{0}(5,\mathbf{p})=(1,\,2,\,1,\,1,\,2)$ and $S_{1}(5,\mathbf{p})=(1,\,1,\,2,\,1,\,2)$. Notice that after the first two seats have been allocated to states 1 and 2 under Adams’ method, the sequence for the next three seats is $1,1,2$, which is the same as the first three seats under d’Hondt’s method. This is because once each state has been awarded a seat, the denominators in each ratio under Adams’ method, $p_{i}/(a_{i}+1),$ are the same as under the d’Hondt method. This relationship holds more generally and leads to the following theorem.
Theorem 1.
For every $\mathbf{p}$, if $S_{1}(\mathbf{p})=(s_{1},s_{2},\dots)$ is the sequential apportionment under D’Hondt’s method then the sequential apportionment under Adams’ method is $S_{0}(\mathbf{p})=(1,2,\dots,n,s_{1},s_{2},s_{3},\dots)$.
Proof.
Under Adams’ method, each of the parties $1,2,\dots,n$ receive, in this order, one of the first $n$ seats. After that, each party has one seat, and so the next seat is given to the party maximizing $\frac{p_{i}}{1}$. These are the same ratios whose maximum determines the first seat under D’Hondt’s method; hence, seat $n+1$ is given to party 1, or to party $s_{1}$ (because $p_{1}\geq p_{i}$ for all $i$). The next seat is given to the party corresponding to the maximum of $\{\{\frac{p_{i}}{1}\}_{i\neq s_{i}},\frac{p_{s_{i}}}{2}\}$. Again, these are the same ratios used to determine the second seat under D’Hondt’s method; so it is given to party $s_{2}$, and so on. ∎
2.1 A Refinement of the Notion of Bias
Many properties of apportionment methods are best described in terms of the number of seats each party receives. Let $F_{c}(h,\mathbf{p})=\mathbf{a}=(a_{1},a_{2},\ldots,a_{n})$ indicate that after $h$ seats, the apportionment method has allocated $a_{i}$ seats to party $p_{i}$, where $h=\sum_{i}a_{i}$. Thus, in Example 2, $F_{0}(5,\mathbf{p})=F_{1}(5,\mathbf{p})=(3,2)$; however $F_{0}(2,\mathbf{p})=(1,1)\neq(2,0)=F_{1}(2,\mathbf{p}).$
When compared to one another, divisor methods are subject to some amount of bias, either favoring large or small parties (based on the number of votes received). The following definition and proposition are due to [2]
Definition 3.
A divisor method $F$ favors large parties relative to divisor method $F^{\prime}$ if for every $\mathbf{p}$ and $h$ with $F(h,\mathbf{p})=\mathbf{a}$ and $F^{\prime}(h,\mathbf{p})=\mathbf{a}^{\prime}$, then $p_{i}>p_{j}$ implies either $a_{i}\geq a_{i}^{\prime}$ or $a_{j}\leq a_{j}^{\prime}$.
Proposition 2.
If $F_{c}$ and $F^{\prime}_{c^{\prime}}$ are stationary divisor methods corresponding to cut points $c$ and $c^{\prime}$, respectively, with $c>c^{\prime}$, then $F$ favors large parties relative to $F^{\prime}$.
Relative biases for stationary divisor methods can also be seen through a lexicographic ordering of their apportionment sequences. For two sequences $s=(s_{1},s_{2},s_{3},\ldots)$ and $t=(t_{1},t_{2},t_{3},\ldots)$, then $s>t$ if $s_{i}=t_{i}$ for $i=1$ to $k$ (so, the sequences agree for the first $k$ places) and $s_{k+1}>t_{k+1}$ for some $k\geq 1$. For a fixed set of votes $p_{i}$ for $i=1$ to $n$, the following proposition shows that apportionment sequences from stationary divisor methods are lexicographically ordered by their cut points.
Proposition 3.
If $c>c^{\prime}$, then $S_{c}(h,\mathbf{p})\leq S_{c^{\prime}}(h,\mathbf{p})$ and $S_{c}(\mathbf{p})\leq S_{c^{\prime}}(\mathbf{p})$ for every $\mathbf{p}$ and $h>0$.
Proof.
Suppose $c>c^{\prime}$ and that there exists $\mathbf{p}$ and an $h$ such that $S_{c}(h,\mathbf{p})>S_{c^{\prime}}(h,\mathbf{p})$. Let $h^{\prime}+1$ be the first index for which $S_{c}(h,\mathbf{p})$ and $S_{c^{\prime}}(h,\mathbf{p})$ differ. Thus
$S_{c}(h^{\prime}+1,\mathbf{p})=(s_{1},s_{2},\ldots,s_{h^{\prime}},i)\quad\mbox{and}\quad S_{c^{\prime}}(h^{\prime}+1,\mathbf{p})=(s_{1},s_{2},\ldots,s_{h^{\prime}},j)$
for some $s_{1},\ldots,s_{h^{\prime}}$ and $i>j$. In addition, $F_{c}(h^{\prime},\mathbf{p})=F_{c^{\prime}}(h^{\prime},\mathbf{p})$; let $\mathbf{a}$ be this common value.
Under $c^{\prime}$, party $j$ gets the $h^{\prime}+1$ seat, which implies
$\frac{p_{j}}{a_{j}+c^{\prime}}>\frac{p_{i}}{a_{i}+c^{\prime}}.$
Under $c$, party $i$ gets the $h^{\prime}+1$ seat, which implies
$\frac{p_{i}}{a_{i}+c}>\frac{p_{j}}{a_{j}+c}.$
Cross multiplying and adding these inequalities yields
$cp_{i}+c^{\prime}p_{j}>cp_{j}+c^{\prime}p_{i}\leftrightarrow(c-c^{\prime})(p_{i}-p_{j})>0.$
Since $c>c^{\prime}$, this implies $p_{i}>p_{j}$ and hence $i<j$. This contradiction proves the claim. ∎
Among stationary divisor methods, D’Hondt’s and Adams’ methods are the most and least biased towards large parties, respectively. Hence, $S_{0}(h,\mathbf{p})\geq S_{1}(h,\mathbf{p})$ for every $h$ and $S_{0}(\mathbf{p})\geq S_{1}(\mathbf{p})$ and, by Proposition 3, $S_{0}(h,\mathbf{p})\geq S_{c}(h,\mathbf{p})\geq S_{1}(h,\mathbf{p})$ for all other cut points $c\in(0,1)$. We will use the lexicographical ordering of apportionment sequences to refine the notion of bias, thereby showing how bias is less an absolute statement and more a reflection of when large or small parties receive their seats.
To achieve this refinement of bias, we need to recall additional properties (homogeneity and weak proportionality) of divisor methods. A divisor method satisfies homogeneity because the apportionment only depends on the relative proportion of votes $p_{i}/P$ where $P=p_{1}+\cdots+p_{n}$. Consequently, for a stationary divisor method with cut point $c$, if $\lambda>0$ is a constant then $F_{c}(h,\mathbf{p})=F_{c}(h,\lambda\mathbf{p})$.
Weak proportionality means that if a vector of seats $\mathbf{a}=(a_{1},a_{2},\ldots,a_{n})$ is proportional to $\mathbf{p}=(p_{1},p_{2},\ldots,p_{n})$, then after $h=\sum_{i}a_{i}$ seats have been allocated, party $i$ has received exactly $a_{i}$ seats, or $F_{c}(\sum_{i}a_{i},\mathbf{p})=\mathbf{a}$. We illustrate this with a continuation of Example 2.
Example 3.
Let $\mathbf{p}=(32,14)$. Since $\mathbf{a}=(16,7)$ and $\mathbf{a}=(32,14)$ are both proportional to $\mathbf{p}$, then under D’Hond’t method, $F_{1}(23,\mathbf{p})=(16,7)$ and $F_{1}(46,\mathbf{p})=(32,14).$ Additionally, their corresponding sequences are
[…]
where $1^{k}$ denotes a sequence of $k$ 1s.
The proof that divisor methods satisfy weak proportionality follows most easily from an equivalent (non-sequential) definition of divisor methods. Under this definition, each apportionment method corresponds to a “rounding rule” $f$ that assigns to each positive integer $k$ a value $f(k)\in[k-1,k]$. Numbers in [k1,f(k))[k-1,f(k)) are rounded to k1k-1 and numbers in $[f(k),k]$ are rounded up to $k$. (The function must also satisfy the property that there cannot exist positive integers $a$ and $b$ for which f(a)=a1f(a)=a-1 and $f(b)=b$.) The divisor method associated with rounding rule $f$ is defined as follows.
Definition 4.
A divisor method $F$ with rounding rule $f$ satisfies $F_{f}(h,\mathbf{p})=\mathbf{a}$ if there exists a nonnegative divisor $d$ such that $f(a_{i})\leq\frac{p_{i}}{d}\leq f(a_{i}+1)$ for each $i$ where $a_{i}$ are nonnegative integers satisfying $\sum_{i=1}^{n}a_{i}=h$.
For stationary divisor methods, f(k)=k1+cf(k)=k-1+c and numbers in [k1,k][k-1,k] are rounded up or down depending on the whether they are above or below k1+ck-1+c (hence the name cut point). For these divisor methods (using the more direct notation $F_{c}$ rather than $F_{f}$), $F_{c}(h,\mathbf{p})=\mathbf{a}$ if there exists a divisor $d$ such that $a_{i}-1+c\leq\frac{p_{i}}{d}\leq a_{i}+c$ for each $i$, or equivalently
$\max_{i}\frac{p_{i}}{a_{i}+c}\leq d\leq\min_{i}\frac{p_{i}}{a_{i}-1+c}.$
The fact that stationary divisor methods can be defined in two ways—through a rounding rule and through a sequential process—is a consequence of the divisor methods’ house monotonicity. An apportionment method is house monotonic if, as the number of seats increases, no party’s allocation decreases. Thus, for stationary divisor methods, and divisor methods more generally, seats can either be allocated simultaneously, by finding an appropriate value of $d$, or sequentially, by increasing the house size one by one. This property is not satisfied by all apportionment methods. (Hamilton’s method, or the method of greatest remainders, is the best-known example of an apportionment method that is not house monotone.)
Divisor methods satisfy weak proportionality since if $\mathbf{p}=k\mathbf{a}$ for some $k\geq 1$, then $\max_{i}\frac{p_{i}}{f(a_{i}+1)}\leq k\leq\min_{i}\frac{p_{i}}{f(a_{i})}.$ If f(k)=k1+cf(k)=k-1+c, this is equivalent to
$\max_{i}\frac{p_{i}}{a_{i}+c}\leq k\leq\min_{i}\frac{p_{i}}{a_{i}-1+c}$
for all $c\in[0,1]$.
2.2 Periodic Behavior
Weak proportionality ensures that an apportionment method allocates seats in exact proportion to the votes received when possible. Clearly this occurs when the house size is a multiple of the total number of votes. It may also occur for smaller values of $h$. In Example 2, the smallest number of house seats $h$ for which an exact proportional allocation is possible is $23$ since $p_{1}=32$ and $p_{2}=14$ are both divisible by $2$ and $(p_{1}+p_{2})/2=23.$
More generally, let $(p_{1}+\cdots+p_{n})/\gcd(p_{1},p_{2},\dots,p_{n})$. It follows that the vote vector $\mathbf{P}=\frac{(p_{1},p_{2},\dots,p_{n})}{\gcd(p_{1},p_{2},\dots,p_{n})}$ is an integer vector that is proportional to $\mathbf{p}$. Thus by both homogeneity and weak proportionality, we have the following proposition.
Proposition 4.
If $F$ is a divisor method and $P=(p_{1}+\cdots+p_{n})/\gcd(p_{1},\ldots,p_{n}),$ then $F(mP,\mathbf{p})=m\cdot\mathbf{P}$ for all integers $m\geq 1$.
From Example 2, we see that not only is $F_{1}(46,\mathbf{p})=2\cdot F_{1}(23,\mathbf{p})$, but the sequence $S_{1}(46,\mathbf{p})$ consists of the sequence $S_{1}(23,\mathbf{p})$ repeated twice. In fact, this periodicity characterizes stationary divisor methods.
Theorem 5.
For a fixed $\mathbf{p}$ and cut point $c$, every stationary divisor method corresponds to a periodic sequential apportionment sequence $S_{c}(\mathbf{p})=s_{1},s_{2},\ldots,$ satisfying $s_{mP+k}=s_{k}$ for all $k=1,\ldots,P$ and $m\geq 1$ where $P=(p_{1}+\cdots+p_{n})/\gcd(p_{1},\dots,p_{n})$.
Proof.
Let $d=\gcd(p_{1},\ldots,p_{n})$ and recall from Proposition 4 that if $\mathbf{P}=\mathbf{p}/d$ then $F_{c}(mP,\mathbf{p})=m\cdot\mathbf{P}$ for all $m\geq 1$. Suppose that $S_{c}(P,\mathbf{p})=(s_{1},\ldots,s_{P})$. We consider which party receives seat $mP+1$. This is awarded to the party maximizing the ratio $\frac{p_{i}}{mP_{i}+c}$ or, equivalently, minimizing the ratios
$\left\{\frac{mP_{i}+c}{p_{i}}\right\}_{i}=\left\{\frac{m}{d}+\frac{c}{p_{i}}\right\}_{i}.$
This is the same party that minimizes $\{\frac{c}{p_{i}}\}_{i}$; hence $s_{mP+1}=s_{1}$.
Seat $mP+2$ seat is awarded to the party that minimizes the ratio
$\left\{\left\{\frac{mP_{i}+c}{p_{i}}\right\}_{i\neq s_{i}},\frac{mP_{s_{1}}+1+c}{p_{s_{1}}}\right\}=\left\{\left\{\frac{m}{d}+\frac{c}{p_{i}}\right\}_{i\neq s_{i}},\frac{m}{d}+\frac{1+c}{p_{s_{1}}}\right\}.$
Again, subtracting the common term $m/d$ this is the same party that received the $2^{nd}$ seat; hence $s_{mP+2}=s_{2}.$
Continuing in this way, we see that $s_{mP+k}=s_{k}$ for $k=1,\ldots,P.$ Hence $S_{c}(\mathbf{p})$ is periodic with period at most $P$.
Now suppose the period is equal to $Q<P$ where $F_{c}(Q,\mathbf{p})=\mathbf{q}$. Then $Q$ divides $P$. Let $\lambda\in\mathbb{N}$ be such an $\lambda Q=P$. By periodicity, $F_{c}(P,\mathbf{p})=F_{c}(\lambda Q,\mathbf{p})=\lambda\cdot F_{c}(Q,\mathbf{p})=\lambda\mathbf{q}$. But $F_{c}(P,\mathbf{p})=\mathbf{P}=\mathbf{p}/d$, which implies $\mathbf{p}/d=\lambda\mathbf{q}$. Thus $\lambda$ divides $p_{i}/d$ for each $i$, which is impossible unless $\lambda=1.$ Hence, $Q=P$. ∎
Although we will usually assume that the $p_{i}$ are natural numbers, so that $p_{i}/p_{j}$ is rational, divisor methods can be equally defined for irrational values. One can imagine contexts in which irrational values could arise in the JIT sequencing context. However, when $p_{i}/p_{j}$ is irrational for some $i$ and $j$, the apportionment sequence is not periodic.
Proposition 6.
If $p_{i}/p_{j}$ is irrational for some $i$ and $j$, then the apportionment sequence is not periodic under any stationary divisor method.
Proof.
Suppose the apportionment sequence is periodic with period $P$ and suppose $F(P,\mathbf{p})=(a_{1},\ldots,a_{n})$ so that $F(mP,\mathbf{p})=(ma_{1},\ldots,ma_{n})$ for all positive integers $m$. Fix $m$ and consider how the $mP+1$ seat would be assigned. If $p_{i}>p_{j}$ then party $i$ receives $ma_{i}+1$ seats before party $j$ receives $ma_{j}+1$ seats, which implies that
[…]
Taking the limit as $m\rightarrow\infty$, $p_{i}/p_{j}\geq a_{i}/a_{j}$.
On the other hand, party $j$ receives $ma_{j}$ seats before party $i$ receives $ma_{i}+1$ seats, leading to the inequalities
[…]
Again, taking the limit as $m\rightarrow\infty$ implies $a_{i}/a_{j}\geq p_{i}/p_{j}$.
Thus the ratio $p_{i}/p_{j}=a_{i}/a_{j}$ is rational, which is a contradiction. Hence, the apportionment sequence cannot be periodic. ∎
Many divisor methods are based on means, in which the cut point between k1k-1 and $k$ is a mean of k1k-1 and $k$. So, the cut point $f(k)$ could be the geometric mean (as in the Hill-Huntington method), the harmonic mean (as in Dean’s method), or any of the parameterized power means. In these cases, the resulting sequence need not be periodic, as demonstrated by the following example. Key here is that all divisor methods satisfy weak proportionality.
Example 4.
Hill-Huntington’s method corresponds to the rounding rule $f(k)=\sqrt{(k-1)k}$. If $\mathbf{p}=(23,4)$ then the sequences for $h=27$ and $h=54$ are
[…]
Notice that the sequence $S_{H}(54,\mathbf{p})$ is not just the sequence $S_{H}(27,\mathbf{p})$ repeated twice. These two sequences of length 54 agree up to the 28th term, but the 29th term in twice-repeated $S_{H}(27,\mathbf{p})$ is a $2$, while the 29th term in $S_{H}(54,\mathbf{p})$ is a 1. Hence, the sequence is not periodic, despite Hill-Huntington’s method being weakly proportional.
In this paper, our primary interest is in understanding the periodic sequences that arise from stationary divisor methods when the $p_{i}$ are rational. In Section 3, we determine explicit expressions for all possible sequences when $n=2$; we extend this analysis to more than two parties in Section 4.
3 Apportionment Sequences for Two Parties
Suppose that there are two parties with vote totals $p_{1}>p_{2}$. Then Theorem 5 states that the sequential apportionment from any stationary divisor is periodic with length $(p_{1}+p_{2})/\gcd(p_{1},p_{2}).$ The exact sequence depends on the value of $c\in[0,1]$.
To understand what these sequences look like, we assume first that $p_{1}$ and $p_{2}$ are relatively prime, so that $\gcd(p_{1},p_{2})=1$. By weak proportionality and homogeneity, each period of length $p_{1}+p_{2}$ consists of $p_{1}$ 1’s and $p_{2}$ 2s. The following lemma places constraints on what sequences are possible.
Lemma 7.
For a 2-party apportionment under a stationary divisor method, if $p_{1}>p_{2}$ party 2 never receives two seats in a row.
Proof.
Fix $0\leq c\leq 1$ and suppose for some $h=a_{1}+a_{2}$, the apportionment is $\mathbf{h}=(a_{1},a_{2})$, and that party 1 receives the next seat followed by 2 seats for party 2. Then
$\frac{p_{1}}{a_{1}-1+c}>\frac{p_{2}}{a_{2}-1+c},\quad\frac{p_{2}}{a_{2}-1+c}>\frac{p_{1}}{a_{1}+c}\quad\mbox{and}\quad\frac{p_{2}}{a_{2}+c}>\frac{p_{1}}{a_{1}+c}.$
The first inequality implies $\frac{(a_{1}-1)p_{2}-(a_{2}-1)p_{1}}{p_{1}-p_{2}}<c$. The third inequality implies $c<\frac{a_{1}p_{2}-a_{2}p_{1}}{p_{1}-p_{2}}$. These two inequalities cannot both be true, so no such $c$ can exist and party 2 cannot get two seats in a row under sequential apportionment. ∎
Combining Lemma 7 and Theorem 5, this means that any two-party apportionment sequence from a stationary divisor method will have a repeated pattern of the form
$S_{c}(p_{1},p_{2})=\underbrace{1,1,\ldots,1}_{k_{1}\text{ times}},2,\underbrace{1,1,\ldots,1}_{k_{2}\text{ times}},2,\ldots,\underbrace{1,1,\ldots,1}_{k_{p_{2}}\text{ times}},2,\underbrace{1,1,\ldots,1}_{p_{1}-\sum_{i}k_{i}\text{ times}}.$
As in Section 2, we use the notation
$S_{c}(p_{1},p_{2})=1^{k_{1}},\,2,\,1^{k_{2}},\,2,\,\cdots\,1^{k_{p_{2}}},\,2,\,1^{p_{1}-(k_{1}+\cdots+k_{p_{2}})},$
where the values of the $k_{i}$ depend on $p_{1}$ and $p_{2}$ as well as on the cut point $c$. This is illustrated in the following example.
Example 5.
Suppose $p_{1}=16$ and $p_{2}=7$. For each $c\in[0,1]$, the apportionment sequence has period $23$. If $c=0$, the first 23 elements of the sequence are
$1,\;2,\;1^{2},\;2,\;1^{2},\;2,\;1^{2},\;2,\;1^{3},\;2,\;1^{2},\;2,\;1^{2},\;2,\;1^{2}.$
For $c$ close to 0, the apportionment sequence remains the same. As $c$ increases in the interval $[0,1]$, there are 10 possible apportionment sequences, arranged in increasing lexicographical order. These are indicated in Table 1. The changes between lines are indicated with the colors red and blue.
Table 1: Apportionment sequences for $p_{1}=16$ and $p_{2}=7$ for stationary divisor methods dependent on $c$.
$c$ Sequence for $p_{1}=16$ and $p_{2}=7$
$\left[0,\frac{1}{9}\right)$ 11 22 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$
$\left[\frac{1}{9},\frac{2}{9}\right)$ 11 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$
$\left[\frac{2}{9},\frac{3}{9}\right)$ 11 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 11
$\left[\frac{3}{9},\frac{4}{9}\right)$ 11 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 11
$\left[\frac{4}{9},\frac{5}{9}\right)$ 11 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 11
$\left[\frac{5}{9},\frac{6}{9}\right)$ 11 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 11
$\left[\frac{6}{9},\frac{7}{9}\right)$ 11 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 11
$\left[\frac{7}{9},\frac{8}{9}\right)$ $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 11
$\left[\frac{8}{9},\frac{9}{9}\right)$ $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 11
11 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22
Note that the apportionment sequences are constant for $c$ in intervals of the form $\left[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}}\right)$ (the half-open intervals are due to the tie-breaking rule), with a final apportionment sequence in which there is no final 1 corresponding to $c=1$. Additionally, $k_{1},k_{2},\ldots,k_{7}\in\{2,3\}.$ This is due to the fact that $2<\frac{23}{7}<3$; hence for the sequence to remain roughly proportional, party 1 must receive 2 or 3 seats for every seat received by party 2.
As a first step towards generalizing these observations, we identify necessary and sufficient conditions for an apportionment sequence to be the result of a stationary divisor method.
Proposition 8.
Suppose $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}>p_{2}$ and the apportionment sequence has the form
$1^{k_{1}}\;2\;1^{k_{2}}\;2\;\ldots\;1^{k_{j}}\;2\;\ldots\;1^{k_{p_{2}}}\;2\;1^{p_{1}-(k_{1}+\cdots+k_{p_{2}})}.$
It follows that
  1. 1.
    The apportionment sequence corresponds to a stationary divisor method if and only if
    $1\leq k_{1}\leq\left\lfloor\frac{p_{1}}{p_{2}}\right\rfloor$ (1)
    and
    $\left\lfloor\frac{(i^{\prime}-i)p_{1}}{p_{2}}\right\rfloor\leq k_{i+1}+\cdots+k_{i^{\prime}}\leq\left\lfloor\frac{(i^{\prime}-i)p_{1}}{p_{2}}\right\rfloor+1\mbox{ for all }1\leq i<i^{\prime}\leq p_{2}.$ (2)
  2. 2.
    If these conditions are met, then the cut point $c$ lies in the interval
    $\left[\max_{i\geq 1}\frac{p_{2}(\sum_{j\leq i}k_{j}-1)-p_{1}(i-1)}{p_{1}-p_{2}},\,\,\min_{i\geq 1}\frac{p_{2}(\sum_{j\leq i}k_{j})-p_{1}(i-1)}{p_{1}-p_{2}}\right).$ (3)
    Moreover, if $C_{c}=\frac{(p_{1}-p_{2})c}{p_{2}}$ then
    $k_{1}$ $=\lfloor C_{c}\rfloor+1,$ (4)
    $k_{i}$ $=\left\lfloor\frac{p_{1}}{p_{2}}(i-1)+C_{c}\right\rfloor-\left\lfloor\frac{p_{1}}{p_{2}}(i-2)+C_{c}\right\rfloor\quad\mbox{ for }i=2,\ldots,p_{2},\text{ and}$
    $k_{p_{2}+1}$ $=p_{1}-\left\lfloor\frac{p_{1}}{p_{2}}(p_{2}-1)+C_{c}\right\rfloor-1.$
We prove Proposition 8 in reverse order.
Proof.
Suppose the apportionment sequence has the indicated form and that it corresponds to a stationary divisor method with cut point $c$. Then for each $i=1,\ldots,p_{2}$, since party 1 receives $k_{1}+\cdots+k_{i}=\sum_{j\leq i}k_{j}$ seats before party 2 receives $i$ seats, $c$ must satisfy (using the tie-breaking rule),
$\frac{p_{1}}{\sum_{j\leq i}k_{j}-1+c}\geq\frac{p_{2}}{i-1+c}>\frac{p_{1}}{\sum_{j\leq i}k_{j}+c}$ (5)
or
$\frac{p_{2}(\sum_{j\leq i}k_{j}-1)-p_{1}(i-1)}{p_{1}-p_{2}}\leq c<\frac{p_{2}(\sum_{j\leq i}k_{j})-p_{1}(i-1)}{p_{1}-p_{2}}.$ (6)
Thus, $c$ must lie in the interval (3).
To show that the $k_{i}$ satisfy (4), we solve for $\sum_{j\leq i}k_{j}$ in (5) to obtain
$\frac{(p_{1}-p_{2})c}{p_{2}}+\frac{p_{1}}{p_{2}}(i-1)<\sum_{j\leq i}k_{j}\leq\frac{(p_{1}-p_{2})c}{p_{2}}+\frac{p_{1}}{p_{2}}(i-1)+1.$
Thus, $\sum_{j\leq i}k_{j}=\lfloor\frac{(p_{1}-p_{2})c}{p_{2}}+\frac{p_{1}}{p_{2}}(i-1)\rfloor+1$ for all $i=1,\ldots,p_{2}$. Substituting $i=1$, $i=2$, etc. sequentially into this expression yields the desired result, and proves the second statement.
The proof of the first statement follows. To show that (1) and (2) are necessary, let $i=1$ in (6) to get
$\frac{p_{2}(k_{1}-1)}{p_{1}-p_{2}}\leq c<\frac{p_{2}k_{1}}{p_{1}-p_{2}}.$
Since $c\in[0,1]$ this implies $\frac{p_{2}(k_{1}-1)}{p_{1}-p_{2}}<1$ and $\frac{p_{2}(k_{1})}{p_{1}-p_{2}}>0$. Solving for $k_{1}$ and recalling that $k_{1}$ is an integer, yields $1\leq k_{1}\leq\left\lfloor\frac{p_{1}}{p_{2}}\right\rfloor$.
To show that (2) is necessary, note that in order for $c$ to exist, the intervals in (6) must have nonempty intersection. Thus, each left end-point must be to the left of each right end-point. So for each $i\neq i^{\prime},$
$\frac{p_{2}(\sum_{j\leq i}k_{j}-1)-p_{1}(i-1)}{p_{1}-p_{2}}<\frac{p_{2}\sum_{j\leq i^{\prime}}k_{j}-p_{1}(i^{\prime}-1)}{p_{1}-p_{2}}.$
If $i<i^{\prime}$, this implies $p_{1}(i^{\prime}-i)<p_{2}(k_{i+1}+\cdots+k_{i^{\prime}}+1)$. If $i^{\prime}<i$ then $p_{1}(i-i^{\prime})>p_{2}(k_{i^{\prime}+1}+\cdots+k_{i}-1)$. Combining these (and switching the labels $i$ and $i^{\prime}$ in the second inequality so that $i^{\prime}>i$) we get
$p_{2}(k_{i+1}+\cdots+k_{i^{\prime}}-1)<p_{1}(i^{\prime}-i)<p_{2}(k_{i+1}+\cdots+k_{i^{\prime}}+1).$
or
$\frac{(i^{\prime}-i)p_{1}}{p_{2}}-1<k_{i+1}+\cdots+k_{i^{\prime}}<\frac{(i^{\prime}-i)p_{1}}{p_{2}}+1.$
Then (2) follows because the $k_{i}$ are integers.
Conversely, suppose that the apportionment sequence has the given form and that (1) and (2) are satisfied. Then the interval in (3) is non-empty and if $c$ lies in the interval, it clearly satisfies (5) for each $i$. Hence the stationary divisor method with cut point $c$ yields the same sequence. It remains to show that the interval (3) lies in $[0,1].$ To see this, note that for each $i\geq 1$,
$p_{2}(k_{1}+\ldots+k_{i})-(i-1)p_{1}$ $\geq p_{2}\left(k_{1}+\left\lfloor(i-1)\frac{p_{1}}{p_{2}}\right\rfloor\right)-(i-1)p_{1}$
$=p_{2}\left(k_{1}+\left\lfloor(i-1)\frac{p_{1}}{p_{2}}\right\rfloor-(i-1)\frac{p_{1}}{p_{2}}\right)$
$\geq p_{2}\left(1+\left\lfloor(i-1)\frac{p_{1}}{p_{2}}\right\rfloor-(i-1)\frac{p_{1}}{p_{2}}\right)>0.$
Thus, the right-hand side of the interval (3) is greater than 0. Similarly, we can show the left-hand side of the interval (3) is less than or equal to 1, implying that $c\in[0,1]$. ∎
We use Proposition 8 to obtain precise expressions for the exponents as follows.
Theorem 9.
Suppose $p_{1}$ and $p_{2}$ are relatively prime and that $p_{1}=ap_{2}+b$ for some integers $a\geq 0$ and $0<b<p_{2}$. If $C_{c}=\frac{(p_{1}-p_{2})c}{p_{2}}=(a-1+\frac{b}{p_{2}})c$ then
$k_{1}$ $=\lfloor C_{c}\rfloor+1$
[…]
$k_{p_{2}+1}$ $=a+b-1+\left\lfloor\frac{b}{p_{2}}(p_{2}-1)+C_{c}\right\rfloor$
where
$S=\left\{i=2,\ldots,p_{2}:i=\left\lceil(j-C_{c}+\left\lfloor C_{c}\right\rfloor)\frac{p_{2}}{b}\right\rceil+1\mbox{, for }j=1,2,\ldots,b\right\}.$
Proof.
The formulas for $k_{1}$ and $k_{p_{2}+1}$ follow directly by substituting $p_{1}=ap_{2}+b$ into (4). Doing the same for the remaining $k_{i}$, we obtain for $i=2,\ldots,p_{2}$,
$k_{i}$ $=\left\lfloor\frac{p_{1}}{p_{2}}(i-1)+C_{c}\right\rfloor-\left\lfloor\frac{p_{1}}{p_{2}}(i-2)+C_{c}\right\rfloor$
$=\left\lfloor a(i-1)+\frac{b}{p_{2}}(i-1)+C_{c}\right\rfloor-\left\lfloor a(i-2)+\frac{b}{p_{2}}(i-2)+C_{c}\right\rfloor$
$=a+\left\lfloor\frac{b}{p_{2}}(i-1)+C_{c}\right\rfloor-\left\lfloor\frac{b}{p_{2}}(i-2)+C_{c}\right\rfloor.$
Since $\frac{b}{p_{2}}<1$, these exponents are equal to either $a$ or $a+1$. We claim at most $b$ of the $k_{i}$ are equal to $a+1$. To see this, we substitute $i^{\prime}=p_{2}$, $i=1$ and $p_{1}=ap_{2}+b$ into (2), to obtain
$\left\lfloor\frac{(p_{2}-1)ap_{2}+b}{p_{2}}\right\rfloor\leq k_{2}+\cdots+k_{p_{2}}\leq\left\lfloor\frac{(p_{2}-1)ap_{2}+b}{p_{2}}\right\rfloor+1,$
which implies $(p_{2}-1)a+\left\lfloor\frac{(p_{2}-2)b}{p_{2}}\right\rfloor\leq k_{2}+\cdots+k_{p_{2}}\leq(p_{2}-1)a+\left\lfloor\frac{(p_{2}-2)b}{p_{2}}\right\rfloor+1.$ Since $(p_{2}-2)b/p_{2}<b$, the claim follows.
Finally, to determine which of the $k_{i}$ is equal to $a+1$, consider the quantity $K_{i}=\lfloor(i-1)\frac{b}{p_{2}}+C_{c}\rfloor$. If $i=1$, $K_{i}=\lfloor C_{c}\rfloor$. As $i$ increases, the values of $K_{i}$ are constant until some smallest $i$ (call it $i_{1}$) satisfies $K_{i_{1}}=\lfloor C_{c}\rfloor+1$; thus $i_{1}$ is the smallest index such that $k_{i}=a+1.$ Similarly, if $i_{2}$ is the smallest $i$ such that $K_{i_{2}}=\lfloor C_{c}\rfloor+2$ then $i_{2}$ is the second smallest index such that $k_{i}=a+1.$ More generally, the indices for which $k_{i}=a+1$ occur at the smallest indices $i_{j}$ for which $K_{i_{j}}=\lfloor C_{c}\rfloor+j.$
Let $\ell_{j}=\lceil(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}\rceil+1$; we claim $i_{j}=\ell_{j}$ for $j\geq 1$. To see this, note that
$(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}\leq\ell_{j}-1<(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}+1,$
which implies
$j+\lfloor C_{c}\rfloor\leq(\ell_{j}-1)\frac{b}{p_{2}}+C_{c}<j+\lfloor C_{c}\rfloor+\frac{b}{p_{2}},$
and hence $\lfloor(\ell_{j}-1)\frac{b}{p_{2}}+C_{c}\rfloor=j+\lfloor C_{c}\rfloor.$
On the other hand
$(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}-1\leq\ell_{j}-2<(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b},$
which implies
$j+\lfloor C_{c}\rfloor-\frac{b}{p_{2}}\leq(\ell_{j}-2)\frac{b}{p_{2}}+C_{c}<j+\lfloor C_{c}\rfloor,$
and so $\lfloor(\ell_{j}-2)\frac{b}{p_{2}}+C_{c}\rfloor<j+\lfloor C_{c}\rfloor.$ Thus, $\ell_{j}$ is the smallest integer such that $K_{\ell_{j}}=\lfloor C_{c}\rfloor+j$ or $i_{j}=\ell_{j}$, as required. ∎
One consequence of Theorem 9 is that if $p_{1}=p_{2}+1$, then the apportionment sequences for all stationary divisor methods with $c\in[0,1)$ coincide.
Proposition 10.
Suppose $p_{1}=p_{2}+1$ for some $p_{2}>1$. Then the apportionment sequence is
[…]
If $p_{2}=1$ and $p_{1}=2$, then the apportionment sequence for all stationary methods is $1^{2},2$.
Proof.
Suppose $p_{1}=p_{2}+1$ for some $p_{2}>1$. Then $C_{c}=(1-1+\frac{1}{p_{2}})c=\frac{c}{p_{2}}$ and $\lfloor C_{c}\rfloor=0$ for all $c\in[0,1]$. Thus, $k_{1}=1$. Moreover, if $c<1$, then (from the definition of the set $S$) we have $\lceil(1-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{1}\rceil+1=\lceil(1-\frac{c}{p_{2}})p_{2}\rceil+1=p_{2}+1$, and so $S$ is empty. Thus, $k_{2}=\cdots=k_{p_{2}}=1$, leading to the apportionment sequence 1, 2, 1, 2,, $1,\,2,\ldots,\,1$. If $c=1$ then $\lceil(1-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{1}\rceil+1=\lceil(1-\frac{c}{p_{2}})p_{2}\rceil+1=p_{2}$, which implies $S=\{p_{2}\}$. Thus $k_{2}=\cdots=k_{p_{2}-1}=1$ and $k_{p_{2}}=2$, leading to the apportionment sequence 1, 2, 1, 2,, $1,\,2,\ldots,\,1^{2},\,2$. The last statement of the proposition follows similarly. ∎
When $p_{1}>p_{2}+1$, we obtain the following.
Theorem 11.
Suppose $p_{1}$ and $p_{2}$ are relatively prime and $p_{1}>p_{2}+1$. Then the apportionment sequences corresponding to stationary divisor methods are constant for cut points in the intervals $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ where $\ell=1,\ldots,p_{1}-p_{2}-1$. Moreover, each of these $p_{1}-p_{2}-1$ sequences is distinct.
Proof.
Suppose $p_{1}=ap_{2}+b$ for some integers $a\geq 0$ and $0<b<p_{2}$ and that $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ for some $\ell=1,\ldots,p_{1}-p_{2}-1$. Let $\ell=\nu p_{2}+\ell^{\prime}$ for some integers $\nu\geq 0$ and $0\leq\ell^{\prime}<p_{2}.$ Then $C_{c}=\frac{p_{1}-p_{2}}{p_{2}}c$ satisfies $\frac{\ell}{p_{2}}\leq C_{c}<\frac{\ell+1}{p_{2}}$, which implies $\lfloor C_{c}\rfloor=\nu$ and $k_{1}=\nu+1$. Thus, $k_{1}$ is constant on this interval.
To show that the other $k_{i}$ are constant on $\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ it suffices to show that the indices $i_{j}=\lceil(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}\rceil+1$ are also constant on this interval. But
$\frac{-\ell^{\prime}-1}{p_{2}}<-C_{c}+\lfloor C_{c}\rfloor\leq\frac{-\ell^{\prime}}{p_{2}},$
which implies
$\frac{jp_{2}-\ell^{\prime}-1}{b}<(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}\leq\frac{jp_{2}-\ell^{\prime}}{b}.$
So $i_{j}=\lceil(j+\lfloor C_{c}\rfloor-C_{c})\frac{p_{2}}{b}\rceil+1=\lceil\frac{jp_{2}-\ell^{\prime}}{b}\rceil+1$ is constant on the interval as claimed.
Finally, to show that each interval of the form $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ corresponds to a different apportionment sequence, it suffices to consider two adjacent intervals, since the sequences increase in lexicographical ordering as $c$ increases.
Suppose that $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ and $c^{\prime}\in[\frac{\ell+1}{p_{1}-p_{2}},\frac{\ell+2}{p_{1}-p_{2}})$ and that the two apportionment sequences are the same. Then $C_{c^{\prime}}=\frac{p_{1}-p_{2}}{p_{2}}c^{\prime}$ satisfies $\frac{\ell+1}{p_{2}}\leq C_{c^{\prime}}<\frac{\ell+2}{p_{2}}$, and recalling that $\ell=\nu p_{2}+\ell^{\prime},$ we obtain $\nu+\frac{\ell^{\prime}+1}{p_{2}}\leq C_{c^{\prime}}<\nu+\frac{\ell^{\prime}+2}{p_{2}}.$ Since the $k_{1}$ values for each sequence are equal, $\lfloor C_{c^{\prime}}\rfloor=\lfloor C_{c}\rfloor=\nu$. Hence $\ell^{\prime}+1<p_{2}$.
In addition, since the $k_{i}$ values for each sequence are equal, we must have for each $j$,
$\left\lceil\frac{jp_{2}-\ell^{\prime}}{b}\right\rceil=\left\lceil(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}\right\rceil=\left\lceil(j-C_{c^{\prime}}+\lfloor C_{c^{\prime}}\rfloor)\frac{p_{2}}{b}\right\rceil=\left\lceil\frac{jp_{2}-(\ell^{\prime}+1)}{b}\right\rceil.$
But $p_{2}$ is relatively prime to $b$. So there exists a $j^{\prime}\in\{1,\ldots,b\}$ such that $j^{\prime}p_{2}$ is congruent to $\ell^{\prime}+1$ mod $b$. Hence $\frac{j^{\prime}p_{2}-(\ell^{\prime}+1)}{b}$ is an integer, which implies $\lceil\frac{j^{\prime}p_{2}-\ell^{\prime}}{p_{2}}\rceil=\lceil\frac{j^{\prime}p_{2}-(\ell^{\prime}+1)}{p_{2}}\rceil+1.$ This contradiction shows that the sequential apportionments corresponding to $c$ and $c^{\prime}$ are distinct. ∎
Since the apportionment is constant on each interval of the form $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$, we can substitute $c=\frac{\ell}{p_{1}-p_{2}}$ into the expressions in Theorem 9 to get the following.
Corollary 12.
Suppose $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}>p_{2}+1$ and $p_{1}=ap_{2}+b$ for some integers $a\geq 0$ and $0<b<p_{2}$. If $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ with $\ell=\nu p_{2}+\ell^{\prime}$ then for $i=2,\ldots,p_{2}$,
$k_{1}$ $=\nu+1,$
$k_{i}$ $=\left\{\begin{array}[]{cl}a+1&\mbox{ if }i=\lceil\frac{jp_{2}-\ell^{\prime}}{b}\rceil+1\mbox{ for some}j=1,2,\ldots,b,\\ a&else,\end{array}\right.\text{ and }$
$k_{p_{2}+1}$ $=p_{1}-(k_{1}+\ldots+k_{p_{2}}).$
Corollary 12 implies that the values of $k_{2},\ldots,k_{p_{2}}$ depend only on $\ell^{\prime}$ (and $b$), and thus are do not change for different values of $k_{1}=\nu+1$. We illustrate this in the following example.
Example 6.
Suppose $p_{2}=7$ and $p_{1}=7a+2$ for some integer $a$. Then for each value of $k_{1}$, there are 7 different apportionment sequences corresponding to $c\in[\frac{\ell}{p_{1}-7},\frac{\ell+1}{p_{1}-7})$ where $\ell=\nu 7+\ell^{\prime}=(k_{1}-1)7+\ell^{\prime}$, as shown in Table 2. For each $\ell^{\prime}=0,\ldots 6$, the position of the exponents for which $k_{i}=a+1$ can be determined by evaluating $i_{j}=\lceil\frac{7j-\ell^{\prime}}{2}\rceil+1.$ Thus, when $\ell^{\prime}=0$ (the first row in the table), this occurs when $i=\lceil\frac{7(1)-0}{2}\rceil+1=5$. When $\ell^{\prime}=1$, this occurs when $i=\lceil\frac{7(1)-1}{2}\rceil+1=4$, and so on.
Table 2: Apportionment sequences for $p_{1}=7a+2$ and $p_{2}=7$.
$c$ Sequences for $p_{1}=7a+2$ and $p_{2}=7$
$\left[\frac{7k_{1}-7}{p_{1}-7},\frac{7k_{1}-6}{p_{1}-7}\right)$ $1^{k_{1}}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a+1}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a-k_{1}-1}$
$\left[\frac{7k_{1}-6}{p_{1}-7},\frac{7k_{1}-5}{p_{1}-7}\right)$ $1^{k_{1}}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a+1}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a-k_{1}-1}$
$\left[\frac{7k_{1}-5}{p_{1}-7},\frac{7k_{1}-4}{p_{1}-7}\right)$ $1^{k_{1}}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a+1}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a+1}$ 22 $1^{a-k_{1}}$
$\left[\frac{7k_{1}-4}{p_{1}-7},\frac{7k_{1}-3}{p_{1}-7}\right)$ $1^{k_{1}}$ 22 $1^{a}$ 22 $1^{a+1}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a+1}$ 22 $1^{a-k_{1}}$
$\left[\frac{7k_{1}-3}{p_{1}-7},\frac{7k_{1}-2}{p_{1}-7}\right)$ $1^{k_{1}}$ 22 $1^{a}$ 22 $1^{a+1}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a+1}$ 22 $1^{a}$ 22 $1^{a-k_{1}}$
$\left[\frac{7k_{1}-2}{p_{1}-7},\frac{7k_{1}-1}{p_{1}-7}\right)$ $1^{k_{1}}$ 22 $1^{a+1}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a+1}$ 22 $1^{a}$ 22 $1^{a-k_{1}}$
$\left[\frac{7k_{1}-1}{p_{1}-7},\frac{7k_{1}}{p_{1}-7}\right)$ $1^{k_{1}}$ 22 $1^{a+1}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a+1}$ 22 $1^{a}$ 22 $1^{a}$ 22 $1^{a-k_{1}}$
When $p_{1}=16$, we retrieve the results in Table 1, where all 7 sequences from Table 2 occur for $k_{1}=1$ and the first 2 sequences from Table 2 occur for $k_{1}=2$.
Table 1 also include a row corresponding to $c=1$ (D’Hondt’s method). It is the only sequence that does not end in a “1.” In fact, this is true more generally.
Proposition 13.
Suppose $p_{1}$ and $p_{2}$ are relatively prime and the apportionment sequence corresponding to the stationary divisor method with cut point $c$ is
$1^{k_{1}}\;2\;1^{k_{2}}\;2\;\ldots 1^{k_{j}}\;2\;\ldots\;1^{k_{p_{2}}}\;2\;1^{p_{1}-(k_{1}+\cdots+k_{p_{2}})}.$
Then $p_{1}=k_{1}+\cdots+k_{p_{2}}$ if and only if $c=1.$
Proof.
Substituting $i=p_{2}$ into (5), we get
$\frac{p_{1}}{\sum_{j\leq p_{2}}k_{j}-1+c}\geq\frac{p_{2}}{p_{2}-1+c}>\frac{p_{1}}{\sum_{j\leq p_{2}}k_{j}+c}.$ (7)
If $c=1$ then $\frac{p_{2}}{p_{2}-1+c}=1,$ which implies
$\frac{p_{1}}{\sum_{j\leq p_{2}}k_{j}}\geq 1>\frac{p_{1}}{\sum_{j\leq p_{2}}k_{j}+1}.$
Hence $\sum_{j\leq p_{2}}k_{j}\leq p_{1}<\sum_{j\leq p_{2}}k_{j}+1,$ which implies $p_{1}=\sum_{j\leq p_{2}}k_{j}.$
Alternatively, if $p_{1}=\sum_{j\leq p_{2}}k_{j}$, then the left-hand inequality in 7) implies
$\frac{p_{1}}{p_{1}-1+c}\geq\frac{p_{2}}{p_{2}-1+c}\Rightarrow c\geq\frac{p_{1}-p_{2}}{p_{1}-p_{2}}=1.$
Note that if $c=1$ and $p_{1}=k_{1}+\cdots+k_{p_{2}}$ then the left-hand inequality in 7 is an equality: $\frac{p_{1}}{p_{1}-1+1}=\frac{p_{2}}{p_{2}-1+1}=1.$ Thus, the last two seats for each party are awarded, first to $p_{1}$ and then to $p_{2}$ using the tie-breaking rule.
Applying Proposition 13 and substituting $c=1$ into the equations in Theorem 9 allows us to arrive at an explicit expression for the apportionment sequence corresponding to D’Hondt’s method.
Proposition 14.
If $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}=ap_{2}+b$, then the apportionment sequence corresponding to D’Hondt’s method is
$1^{a},\,2,\,1^{k_{2}},\,2,\,1^{k_{3}},\,2,\,\ldots,1^{k_{p_{2}}},\,2$
where $k_{i}=a+1$ if $i\in S$ and $k_{i}=a$ otherwise, and $S=\{\lceil\frac{p_{2}}{b}\rceil,\lceil\frac{2p_{2}}{b}\rceil,\ldots,\lceil p_{2}\rceil\}.$
Since $c=1$ corresponds to a different apportionment sequence from any $c\in[0,1)$, Proposition 11 implies the following.
Corollary 15.
If $p_{1}$ and $p_{2}$ are relatively prime then there are exactly $p_{1}-p_{2}+1$ different apportionment sequences corresponding to stationary divisor methods.
Finally, we note we can also use the expressions in Corollary 12 to work backwards from an apportionment sequence to find its corresponding cut point. The following Theorem, whose proof follows directly from solving for $\nu$ and $\ell$ in Corollary 12, provides an explicit formula for the interval that $c$ must lie in given an apportionment sequence with the appropriate form.
Theorem 16.
Suppose $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}>p_{2}+1$ and the apportionment sequence has the form
$1^{k_{1}},\,2,\,1^{k_{2}},\,2,\,\ldots,1^{k_{p_{2}}},\,2,\,1^{p_{1}-(k_{1}+\cdots+k_{p_{2}})}.$
where the $k_{i}$ satisfy (1) and (2). Then the apportionment sequence corresponds to a stationary divisor method with cut point $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}+p_{2}})$ where $\ell=\nu p_{2}+\ell^{\prime}$, $\nu=k_{1}-1$ and $\ell^{\prime}=\max_{i_{j}\in S}\{p_{2}j-(i_{j}-1)b\}$ where $S=\{i_{j},|k_{i_{j}}=\lfloor\frac{p_{1}}{p_{2}}\rfloor+1,2\leq i_{j}\leq p_{2}\}$.
We conclude this by remarking that all the results of this section apply to more general values of $\mathbf{p}=(p_{1},p_{2})$ with $p_{1}$ and $p_{2}$ replaced by $p_{1}/\gcd(p_{1},p_{2})$ and $p_{2}/\gcd(p_{1},p_{2})$ in all expressions for $k_{i}$, etc.
4 Apportionment Sequences for More Than Two Parties
The properties of apportionment sequences for two parties carry over to apportionment sequences for more than two parties because of the property of consistency. Consistency requires that the apportionment between two states is unaffected by the addition of additional states. More precisely, suppose $h$ seats are allocated among $n$ parties according to $F(h,\mathbf{p})=\mathbf{a}$ and suppose $N^{\prime}$ is a subset of parties whose total number of seats in this allocation is $\sum_{i\in N^{\prime}}a_{i}=h^{\prime}$. If we remove these seats and reapportion the remaining $h-h^{\prime}$ seats among the parties not in $N^{\prime}$, their allocation is unchanged.
All divisor methods satisfy consistency because they are based on ratios of the form $p_{i}/p_{j}$ which are unchanged if additional parties are added or eliminated [2]. From a sequential apportionment point of view, consistency implies that given an apportionment sequence among $n$ parties, if we extract only the entries corresponding to parties $i$ and $j$, the sequence of $i$’s and $j$’s will be the same as the sequence of $1$s and $2$s generated by apportioning seats only between those two parties.
Because of this, some of the results for 2-party sequences extend easily to more than two parties. Lemma 7 is generalized in the following proposition.
Proposition 17.
For an $n$-party apportionment sequence under a stationary divisor method, if $p_{k}>p_{\ell}$ and party $k$ receives a seat in the sequence, then party $\ell$ cannot receive two seats before party $k$ receives another seat.
Proof.
For any divisor method, the subsequence with $i$ and $i$ as its terms is the same as the 2-party sequence when $i$ is identified with 1 and $j$ is identified with 2. By Lemma 7, the sequence with 1s and 2s cannot have two 2s in a row. Thus, two $j$s cannot appear between consecutive $i$s. ∎
In what follows, we extend the results from Section 3 on 2-party apportionment sequences and explain how to construct $n$-party sequences from all $\binom{n}{2}$ 2-party sequences. To demonstrate the purely mechanical process, we first consider an example. After the example, we explain why this process is valid in general.
Example 7.
Let $p_{1}=16$, $p_{2}=11$, and $p_{3}=7$ and $c\in[0,\frac{1}{9})$. Applying Proposition 11, we list the the pairwise sequential apportionment sequences for $p_{1},p_{2}$ and $c\in[0,1/5)$ (first row in Table 3); for $p_{1},p_{3}$ and $c\in[0,1/9)$ (second row in Table 3); and for $p_{2},p_{3}$ and $c\in[0,1/4)$ (third row in Table 3). We will combine the 2-party sequences to form a periodic 3-party sequence of period $h=p_{1}+p_{2}+p_{3}$ for $c\in[0,\frac{1}{9})$, the intersection of $0,\frac{1}{5})$, $0,\frac{1}{9})$, and $0,\frac{1}{4})$. Notice that rows 1, 2, and 3 of Table 3 are of length 27, 23, and 18, because each pair of $p_{i}$ and $p_{j}$ are relatively prime and the sequences are periodic of length $p_{i}+p_{j}$.
Table 3: The three rows in this table are the apportionment sequences for parties 1 and 2 for $c\in[0,\frac{1}{5})$, for parties 1 and 3 for $c\in[0,\frac{1}{9})$, and for parties 2 and 3 for $c\in[0,\frac{1}{4})$, respectively.
11 22 11 22 11 22 11 11 22 11 22 11 11 22 11 22 11 11 22 11 22 11 11 22 11 22 11
11 33 11 11 33 11 11 33 11 11 33 11 11 11 33 11 11 33 11 11 33 11 11
22 33 22 33 22 22 33 22 33 22 22 33 22 33 22 22 33 22
To construct the 3-party sequence, we examine the three 2-party sequences from left to right in Table 3. The top two sequences each begin with a 1. Hence, the 3-party sequence begins with 1. Now, imagine crossing out the 1s. The leading left (non-crossed out) entries in rows 1, 2, and 3 are now 2, 3, and 2, respectively. Two of the rows match with 2, and 2 is the next term in the 3-party sequence. To determine the third term in the 3-party sequence, we notice that the non-used leading left entries are now 1, 3, and 3, which implies that the third term in the sequence is 3. Continuing in this fashion gives the 3-party apportionment sequence of period $34$:
11 22 3 11 22 11 3 22 11 11 22 3 11 22 11 3 11 22 11 22 11 3 11 22 11 33 22 11 11 22 33 11 22 11.
We refer to the process described in Example 7 as lifting. More generally, given $n$ parties with votes $p_{i}$ for $i=1$ to $n$ and cut point $c$, we can define an algorithm to lift the $\binom{n}{2}$ 2-party apportionment sequences corresponding to $c$ to an $n$-party apportionment sequence for $c$. The algorithm begins by generating the $\binom{n}{2}$ 2-party apportionment sequences. The key step is that there is always a unique number between 1 and $n$ that begins n1n-1 of the sequences. This is the next seat to be awarded in the $n$-party apportionment sequence. After awarding this seat, that number is removed from its leftmost position in the n1n-1 sequences and the process is repeated.
Algorithm 1 Lifting from all $\binom{n}{2}$ 2-party sequences to the $n$-party sequence
Input: $n,c,$ and $p_{i}$ for $i=1$ to $n$
Let $m=\gcd(p_{1},p_{2},\dots,p_{n})$
Let $p=(p_{1}+p_{2}+\cdots+p_{n})/m$
Output: combinatorial word $w=\{w_{i}\}_{i=1}^{p}$ where $w_{i}\in\{1,2,\dots,n\}$
for $i=1$ to $n$ do
  Let $j=i+1$
  while $j\leq n$ do
   Let $\ell=(p_{i}+p_{j})/m$
   Generate the 2-party sequence for parties $i$ and $j$ of length $\ell$ (using $i$ and $j$)
   Return line
   $j=j+1$
  end while
end for
for $i=1$ to $p$ do
  for $j=1$ to $\frac{n(n+1)}{2}$ do
   for $k=1$ to $n$ do
     $n_{k}=0$
   end for
   if the leftmost entry in the $j$th 2-party sequence is $k$ then
     $n_{k}=n_{k}+1$
   end if
  end for
  for $k=1$ to $n$ do
   if $n_{k}=n-1$ then
     $w_{i}=k$
   end if
  end for
  Delete the leftmost entry of the rows that begin with $w_{i}$
end for
The following theorem proves that the algorithm generates the $n$-party apportionment sequence. In particular, the proof explains why there is always a number that appears n1n-1 times as the leftmost entry of the 2-party sequences and that the sequence terminates with all 2-party sequences’ entries being emptied.
Theorem 18.
Algorithm 1 generates the $n$-party apportionment sequence from all $\binom{n}{2}$ 2-party sequences.
Proof.
First, we show that at every step of the algorithm there are n1n-1 leftmost entries that are the same. Assume that $k$ seats have been allocated in the $n$-party apportionment sequence and that each party $i$ receives $a_{i}$ seats. The party that receives seat $k+1$ is the party $\ell$ that satisfies $\arg\max_{i\in N}\frac{p_{i}}{a_{i}-1+c}$. If more than one party satisfies the condition, then we apply the tie-break rule that smallest party number receives the seat. There are n1n-1 2-party apportionment sequences that have $\ell$ as one of the two parties. Because $\frac{p_{\ell}}{a_{\ell}-1+c}\geq\frac{p_{i}}{a_{i}-1+c}$ for all $i$, then $\ell$ appears as the leftmost entry in those n1n-1 sequences. Hence, there is always a party that is leftmost in n1n-1 sequences, as required in Algorithm 1. This party is clearly unique since in the 2-sequence between $\ell$ and any other party, $i$, party $i$ will not appear next.
Second, we show that $\frac{p_{1}+p_{2}+\cdots+p_{n}}{\gcd(p_{1},p_{2},\ldots,p_{n})}$ seats are allocated and that no extra terms are left in the 2-party sequences. Each time a seat is awarded, then the (same) leftmost characters are removed n1n-1 2-party sequences. For parties $i$ and $j$, the 2-party sequence from the algorithm consists of $\frac{p_{i}+p_{j}}{\gcd(p_{1},p_{2},\ldots,p_{n})}$ terms. Hence, there are a total of
$\sum_{i\neq j}\frac{p_{i}+p_{j}}{\gcd(p_{1},p_{2},\ldots,p_{n})}=(n-1)\cdot\frac{p_{1}+p_{2}+\cdots+p_{n}}{\gcd(p_{1},p_{2},\ldots,p_{n})}$
terms. Since each step of the algorithm removes n1n-1 terms, there are $\frac{p_{1}+p_{2}+\cdots+p_{n}}{\gcd(p_{1},p_{2},\ldots,p_{n})}$ seats awarded as required. ∎
Example 8.
To determine all possible sequential apportionments for $p_{1}=16$, $p_{2}=11$, and $p_{3}=7$, we can apply Theorem 18 for all $c$ by comparing the pairwise sequential apportionments for $p_{1},p_{2}$ (Table 4), for $p_{1},p_{3}$ (Table 1 with the 2s replaced by 3s), and for $p_{2},p_{3}$ (Table 4). The combined apportionments appear in Table 5; this table includes colored entries to highlight how the sequences changes as the value of $c$ increases.
Table 4: The pairwise sequential apportionments for $p_{1}=16,p_{2}=11$ and $p_{2}=11,p_{3}=7$.
$c$ Sequence for $p_{1}=16$ and $p_{2}=11$
$\left[0,\frac{1}{5}\right)$ 11 22 11 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 11
$\left[\frac{1}{5},\frac{2}{5}\right)$ 11 22 11 22 $1^{2}$ 22 11 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 11
$\left[\frac{2}{5},\frac{3}{5}\right)$ 11 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 11
$\left[\frac{3}{5},\frac{4}{5}\right)$ 11 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 11 22 $1^{2}$ 22 11 22 11
$\left[\frac{4}{5},\frac{5}{5}\right)$ 11 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 11 22 11
11 11 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22 11 22 $1^{2}$ 22
$c$ Sequence for $p_{2}=11$ and $p_{3}=7$
$\left[0,\frac{1}{4}\right)$ 22 33 22 33 $2^{2}$ 33 22 33 $2^{2}$ 33 22 33 $2^{2}$ 33 22
$\left[\frac{1}{4},\frac{2}{4}\right)$ 22 33 22 33 $2^{2}$ 33 22 33 $2^{2}$ 33 $2^{2}$ 33 22 33 22
$\left[\frac{2}{4},\frac{3}{4}\right)$ 22 33 22 33 $2^{2}$ 33 $2^{2}$ 33 22 33 $2^{2}$ 33 22 33 22
$\left[\frac{3}{4},\frac{4}{4}\right)$ 22 33 $2^{2}$ 33 22 33 $2^{2}$ 33 22 33 $2^{2}$ 33 22 33 22
11 22 33 $2^{2}$ 33 22 33 $2^{2}$ 33 22 33 $2^{2}$ 33 $2^{2}$ 33
Table 5: The sequential apportionments for $p_{1}=16$, $p_{2}=11$, and $p_{3}=7$.
$c$ Sequence for $p_{1}=16$, $p_{2}=11$, and $p_{3}=7$
$0,\frac{1}{9})$ 1 2 3 1 2 1 3 2 1 1 2 3 1 2 1 3 1 2 1 2 1 3 1 2 1 3 2 1 1 2 3 1 2 1
$\frac{1}{9},\frac{1}{5})$ 1 2 3 1 2 1 3 2 1 1 2 3 1 2 1 1 3 2 1 2 1 3 1 2 1 3 2 1 1 2 3 1 2 1
$\frac{1}{5},\frac{2}{9})$ 1 2 3 1 2 1 3 1 2 1 2 3 1 2 1 1 3 2 1 2 1 3 1 2 1 3 2 1 1 2 3 1 2 1
$\frac{2}{9},\frac{1}{4})$ 1 2 3 1 2 1 3 1 2 1 2 3 1 2 1 1 3 2 1 2 1 3 1 2 1 3 2 1 1 2 1 3 2 1
$\frac{1}{4},\frac{3}{9})$ 1 2 3 1 2 1 3 1 2 1 2 3 1 2 1 1 3 2 1 2 1 3 1 2 1 2 3 1 1 2 1 3 2 1
$\frac{3}{9},\frac{2}{5})$ 1 2 3 1 2 1 3 1 2 1 2 1 3 2 1 1 3 2 1 2 1 3 1 2 1 2 3 1 1 2 1 3 2 1
$\frac{2}{5},\frac{4}{9})$ 1 2 3 1 2 1 3 1 2 1 2 1 3 1 2 1 3 2 1 2 1 3 1 2 1 2 3 1 1 2 1 3 2 1
$\frac{4}{9},\frac{2}{4})$ 1 2 3 1 2 1 3 1 2 1 2 1 3 1 2 1 3 2 1 2 1 3 1 2 1 2 1 3 1 2 1 3 2 1
$\frac{2}{4},\frac{5}{9})$ 1 2 3 1 2 1 3 1 2 1 2 1 3 1 2 1 2 3 1 2 1 3 1 2 1 2 1 3 1 2 1 3 2 1
$\frac{5}{9},\frac{3}{5})$ 1 2 3 1 2 1 1 3 2 1 2 1 3 1 2 1 2 3 1 2 1 3 1 2 1 2 1 3 1 2 1 3 2 1
$\frac{3}{5},\frac{6}{9})$ 1 2 3 1 2 1 1 3 2 1 2 1 3 1 2 1 2 3 1 1 2 3 1 2 1 2 1 3 1 2 1 3 2 1
$\frac{6}{9},\frac{3}{4})$ 1 2 3 1 2 1 1 3 2 1 2 1 3 1 2 1 2 3 1 1 2 1 3 2 1 2 1 3 1 2 1 3 2 1
$\frac{3}{4},\frac{7}{9})$ 1 2 3 1 2 1 1 2 3 1 2 1 3 1 2 1 2 3 1 1 2 1 3 2 1 2 1 3 1 2 1 3 2 1
$\frac{7}{9},\frac{4}{5})$ 1 2 1 3 2 1 1 2 3 1 2 1 3 1 2 1 2 3 1 1 2 1 3 2 1 2 1 3 1 2 1 3 2 1
$\frac{4}{5},\frac{8}{9})$ 1 2 1 3 2 1 1 2 3 1 2 1 3 1 2 1 2 3 1 1 2 1 3 2 1 1 2 3 1 2 1 3 2 1
$\frac{8}{9},1)$ 1 2 1 3 2 1 1 2 3 1 2 1 3 1 2 1 2 1 3 1 2 1 3 2 1 1 2 3 1 2 1 3 2 1
11 1 2 1 3 2 1 1 2 3 1 2 1 3 1 2 1 2 1 3 1 2 1 3 2 1 1 2 3 1 2 1 1 2 3
Consistency also allows us to characterize all apportionment sequences for $n$ parties corresponding to stationary divisor methods.
Theorem 19.
Given an apportionment sequence of $n$ parties $S=s_{1},s_{2},s_{3},\ldots$, let $S_{i,j}$ be the subsequence created by extracting all the $i$’s and $j$s, for all $i<j$, and let $S^{\prime}_{i,j}$ correspond to $S_{i,j}$ with $i$ and $j$ replaced by $1$ and $2$ respectively. Then $S$ corresponds to a stationary divisor method if and only if $S_{i,j}^{\prime}$ satisfies equations (1) and (2) from Proposition 8 for all $i<j$ and if the intervals (3) have nonempty intersection. If that is the case, then $S$ corresponds to the stationary divisor method for all $c$ in that intersection.
4.1 Counting the number of distinct apportionment sequences for $n$ parties
We can also determine the number of distinct apportionment sequences for $n$ parties. Recall from Section 3, if $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}>p_{2}$, the interval $[0,1)$ can be partitioned into $p_{1}-p_{2}$ intervals of the form $\frac{\ell}{p_{1}-p_{2}},\frac{\ell}{p_{1}-p_{2}})$. Let $A=\{0,\frac{1}{p_{1}-p_{2}},\frac{2}{p_{1}-p_{2}},\ldots,\frac{p_{1}-p_{2}-1}{p_{1}-p_{2}}\}$ be the set of endpoints of these intervals (not including 1). Then the number of apportionment sequences is equal to $|A|+1$, where the extra $1$ corresponds to $c=1$. If $p_{1}$ and $p_{2}$ are not relatively prime, the same thing is true, where $p_{1}$ and $p_{2}$ are replaced with $p_{1}/\gcd(p_{1},p_{2})$ and $p_{2}/\gcd(p_{1},p_{2}).$
Now suppose there are $n$ parties. Each of the $\binom{n}{2}$ pairs generates its own partition of the interval $[0,1)$ and its own set of endpoints $A$. When we combine these pairs to enumerate the sequences for all $n$ parties, as in Example 8, we obtain a refinement of each of the partitions of $[0,1)$ based on the union of the sets $A$.
To determine the endpoints in this union, we require the following Lemma, which follows easily from induction.
Lemma 20.
Let $A_{i}=\{0,\frac{1}{m_{i}},\ldots,\frac{m_{i}-1}{m_{i}}\}$ for some integers $m_{i}$, $i=1,\ldots,M$. Then for all $1\leq i_{1}<i_{2}<\cdots<i_{k}\leq M,$
$A_{i_{1}}\cap\cdots\cap A_{i_{k}}=\left\{0,\frac{1}{m},\ldots,\frac{m-1}{m}\right\},$
where $m=\gcd(m_{i_{1}},\ldots,m_{i_{k}}).$
Counting the number of possible sequential apportionments for $c\in[0,1)$ is then an application of the principle of inclusion-exclusion. To make the inclusion-exclusion cleaner, let $T=\{1,2,\ldots,\binom{n}{2}\}$ be an enumeration of all possible pairs of states (the order is irrelevant). For each $k\in T$, if $k$ corresponds to $p_{i}$ and $p_{j}$ with $p_{i}>p_{j}$, then let $m_{k}=(p_{i}-p_{j})/\gcd(p_{i},p_{j})$ and let $A_{k}=\{0,1/m_{k},2/m_{k},\ldots,(m_{k}-1)/m_{k}\}$. We assume first that each of the $p_{i}$ is distinct.
Theorem 21.
Let $p_{1},p_{2},\dots,p_{n}$ be the vote totals of $n$ parties where $p_{1}>p_{2}\cdots>p_{n}$. The number of possible sequential apportionments is
$1+\left|\bigcup_{k=1}^{\binom{n}{2}}A_{k}\right|$ $=1+\sum_{k=1}^{\binom{n}{2}}(-1)^{k+1}\left(\sum_{1\leq i_{1}<\cdots<i_{k}\leq\binom{n}{2}}|A_{i_{1}}\cap\cdots\cap A_{i_{k}}|\right)$
$=1+\sum_{k=1}^{\binom{n}{2}}(-1)^{k+1}\left(\sum_{1\leq i_{1}<\cdots<i_{k}\leq\binom{n}{2}}\gcd(m_{i_{1}},\ldots,m_{i_{k}})\right).$
Example 9.
Let $\mathbf{p}=(25,17,13,5)$. Then,
$m_{1}=8\quad m_{2}=12\quad m_{3}=4\quad m_{4}=4\quad m_{5}=12\quad m_{6}=8.$
Applying Theorem 21 and not reducing fractions so it is clear where the term originates, $\left|\cup_{k=1}^{6}A_{k}\right|=48-72+80-60+24-4=16$ from the refined partition of $[0,1)$ given by $\cup_{k=1}^{6}A_{k}$ which is equal to
$\{0,1/12,1/8,2/12,1/4,4/12,3/8,5/12,2/4,7/12,5/8,8/12,3/4,10/12,7/8,11/12\}.$
Thus, there are $17$ possible sequential apportionments under stationary divisor methods.
If the party votes are not distinct, Theorem 21 can easily be modified. Note that if two parties have equal votes (say $p_{i}=p_{i+1}$), then the addition of party $i+1$ does not introduce any new endpoints to the 2-party apportionment sequences with parties with non-equal vote totals. Additionally, whenever one of the parties with equal vote totals is awarded a seat (the one with the least index by the tie-breaking rule), then the next largest indexed party with the same vote total gets the next seat, and this continues until all parties with equal vote totals get a seat. This leads to the following corollary.
Corollary 22.
(to Theorem 21). Let $p_{1},p_{2},\dots,p_{n}$ be the vote totals of parties with $p_{i}\geq p_{j}$ for $i<j$. Let $N^{*}$ be one of the largest subsets of $N$ with all $p_{i}$ distinct. Relabel the elements in $N^{*}$ as $q_{1},q_{2},\ldots,q_{m}$ with $q_{i}>q_{j}$ for $i<j$. Then, then the number of possible $n$-party apportionment sequences of the parties in $N$ is the number of possible $m$-party apportionment sequences of the parties in $N^{*}$.
5 Conclusion
The selection of the Northern Ireland Assembly Executive Committee was based on several pre-determined agreements as well as the sequential use of D’Hondt’s method. Nonetheless, it provides a good example with which to consider the impact of sequences drawn from other apportionment methods. For instance, if we apply Theorem 21 to the top four parties who received seats, SF, DUP, AP and UUP with vote distribution $\mathbf{p}=(27,25,17,9)$, there are a total of 25 possible different sequences arising from stationary apportionments. Of course, these have length $27+25+17+9=78$. If we are interested only in the first 10 positions, we can compare sequences directly. Assuming that all parties start with no seats automatically granted, under D’Hond’ts method, the first 10 terms of the sequence are: $1,2,3,1,2,1,4,3,2,1$ (where we have used the tie break to award seat 6 to party 1).
Under Adams’ method, (applying Theorem 1), the first 10 terms of the sequence are: $1,2,3,4,1,2,3,1,2,1$, where the tie is broken in favor of SF in awarding the last seat. Both of these methods begin $1,2,3$ and both award the four parties $4,3,2$ and $1$ seats respectively (as occurred in real life); hence this will be true of all divisor methods. However, the two sequences would have very different political consequences if used to select executive committee positions one at a time. Neither of these sequences match what actually occurred, which was equivalent to the sequence $1,2,3,1,2,1,2,4,1,3$, although of course, the first three positions of First Minister, Vice Minister, and Justice Minister match what the apportionment sequences suggest. It would be interesting to investigate the consequences of different apportionment sequences in other political contexts.
Awarding representative seats sequentially in a consistent manner is a consequence of an apportionment method satisfying house monotonicity. Although apportionment methods have been applied to award cabinet minister positions sequentially and to just-in-time sequencing problems, the study of what is possible under sequential apportionment is underdeveloped. This work is a first in trying to understand both properties of $n$-apportionment sequences as well as the number of possible sequences generated by stationary divisor methods. Because divisor methods are weakly proportional, every party will eventually receive a number of seats equal to its vote totals if the house size is the sum of the votes. This motivates an alternate way of viewing size bias to include when a parties receives its representatives, as opposed to how many. This has lead to a partial order on sequences bookended by the sequences from the methods of Adams and d’Hondt, which favor, respectively, the smallest and largest parties.
For 2-party sequences, we were able to understand all possible sequences. And, by using the 2-party sequences, we were able to determine how many possible $n$-party apportionment sequences there are. There is still room for additional work in this area. By understanding what sequences are possible, it may entice practitioners to develop additional applications of sequential apportionment besides the awarding of cabinet minister positions and its use in just-in-time sequencing.